OpenStudy (anonymous):

Helium is pumped into a spherical balloon at a rate of 3 cubic feet per second. How fast is the radius increasing after 2 minutes? Note: The volume of a sphere is given by V=(4/3)πr^3

OpenStudy (dumbcow):

differentiate volume equation wrt "t" \[\frac{dV}{dt} = 4 \pi r^2 \frac{dr}{dt}\] dV/dt = 3 if initial volume is 0 then after 2 min (120 sec) the new volume is 360 solve for radius , plug into derivative equation to find dr/dt